OpenStudy (anonymous):

log(x+16)-log(x)=1 solve for x

jhonyy9 (jhonyy9):

so do you know the property of logarithms ?

jhonyy9 (jhonyy9):

lod a - log b = ?

OpenStudy (anonymous):

log(x(x+16)=1?

jhonyy9 (jhonyy9):

not is right because log a -log b =log a/b

OpenStudy (anonymous):

oh okay so log(x+16)/(x)=1?

jhonyy9 (jhonyy9):

yes can you tel me in this your exersice what is the base of log x ?

OpenStudy (anonymous):

10

jhonyy9 (jhonyy9):

this is sure ?

OpenStudy (anonymous):

yes sir

jhonyy9 (jhonyy9):

ok,nice so than the 1 how can you rewriting using logarithm

jhonyy9 (jhonyy9):

suppose you know it that log(a) a =1

jhonyy9 (jhonyy9):

came on this is easy

OpenStudy (anonymous):

hm I'm not sure... are you refering to log(x+16)(x+16) ?

jhonyy9 (jhonyy9):

so because you know that log x has a base of 10 so than you can rewriting the 1 using logarithm with base of 10 like 1=log 10

jhonyy9 (jhonyy9):

ok ?

OpenStudy (anonymous):

oooooh

jhonyy9 (jhonyy9):

yes ?

jhonyy9 (jhonyy9):

so than how will be your exersice using this ?

OpenStudy (anonymous):

so you get log(x+16)/(x)=log10 ?

OpenStudy (anonymous):

but I dont know what the step is after that..

jhonyy9 (jhonyy9):

yes

jhonyy9 (jhonyy9):

so when will be equal log a =log b ?

jhonyy9 (jhonyy9):

how do you think log 2 =log 3 ?

jhonyy9 (jhonyy9):

or log 5=log 5 ?

OpenStudy (anonymous):

its multiplied by 1?

jhonyy9 (jhonyy9):

no when will be equal log a = log b ?

OpenStudy (anonymous):

Im sorry i just dont know

jhonyy9 (jhonyy9):

so you know that log 5 = log 5 because 5=5 ok ? so this mean that log a =log b just when a=b this mean in your case what ?

OpenStudy (anonymous):

x+16=x?

jhonyy9 (jhonyy9):

no check it please

OpenStudy (anonymous):

x+16=-x?

jhonyy9 (jhonyy9):

the left part equal right part so (x+16)/x =10 right ?

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

why ?

OpenStudy (anonymous):

because the log cancel out?

jhonyy9 (jhonyy9):

ok can you continue it to solve for x ?

jhonyy9 (jhonyy9):

(x+16)/x =10 what is the next step ?

OpenStudy (anonymous):

well i multiplied X by both sides and that gave me x+16=10x

jhonyy9 (jhonyy9):

yes and ?

OpenStudy (anonymous):

well i got x=1.78

jhonyy9 (jhonyy9):

how ?

OpenStudy (anonymous):

16=9x 1.78=x x=1.78

jhonyy9 (jhonyy9):

yes right good luck was my pleasure bye

OpenStudy (anonymous):

um well i plugged that answer into my computer program (Wiley Plus) and it said that was the incorrect answer. are you sure i did it correctly?

OpenStudy (accessdenied):

Are you sure you typed it correctly? Otherwise, could it be looking for an exact figure: 16/9? The answer should be correct.